# Model Collapse in the Self-Consuming Chain of Diffusion Finetuning: A Novel Perspective from Quantitative Trait Modeling

### v2

@@ 4 Quantitative Trait Models for Fully-Synthetic Training Loops
This section introduces a novel perspective for understanding the model collapse in generative models by drawing parallels between genetic biology and the Chain of Diffusion. Distinguishable iterations in the Chain of Diffusion—where each iteration is clearly separable with no duplicated individuals, and the current iteration originates from the previous one—mirror genetic processes involving successive iterations of parents and offspring. By applying quantitative trait modeling from statistical genetics, we provide a framework to describe how images evolve across iterations in the Chain of Diffusion.
We begin by introducing quantitative trait modeling and its underlying mathematical assumptions. Leveraging these assumptions, we derive a theorem demonstrating that the mean trait value exhibits linear drift while the variance stabilizes over time. We then show that this model can successfully capture the three key behaviors observed in Section 3: high-frequency degradation, low-frequency degradation, and optimal CFG scale performance. This suggests that different CFG scales in the Chain of Diffusion can be viewed as varying selection strategies, with power spectra corresponding to quantitative traits. We use the term “iteration” instead of “generation” for genetic generation to avoid confusion with image generation.
(a) Selection mechanism.
(b) Power spectra of images vs. Mean of phenotypes.
Figure 4: (a) Selection mechanism with two-sided truncation. $r_{1}$ and $r_{2}$ ratios of samples are truncated from the left and right tails, and the remaining $r=1-r_{1}-r_{2}$ ratio of samples is selected. (b) Power spectra for different CFGs (Pokemon) and quantitative trait modeling results of different selection strategies. Directional selections with truncation can effectively explain our observations in Section 3: the behaviors of high- and low-frequency degradation and optimal CFG scale. Detailed settings are provided in Appendix D.2.
@@ 4.1 Quantitative Trait Modeling
Quantitative trait modeling in statistical genetics explores the evolution of quantitative phenotypes (e.g., height, weight, or color), which are determined by multiple genetic and environmental factors in complex ways. These phenotypes are typically assumed to follow a Gaussian distribution, with discrete iterations where parents and offspring are distinguishable ($t$- and $t+1$-th iterations are clearly separable).
Let the distribution of phenotypes at the $t$-th iteration be denoted as $\mathcal{N}(\mu_{t},\sigma_{P,t}^{2})$ where $\mu_{t}$ and $\sigma_{P,t}^{2}$ are the mean and variance of quantitative phenotypes. The phenotypic variance $\sigma_{P,t}^{2}$ is the sum of the (additive) genetic variance $\sigma_{G,t}^{2}$ and the environmental variance $\sigma_{E}^{2}$, i.e., $\sigma_{P,t}^{2}=\sigma_{G,t}^{2}+\sigma_{E}^{2}$ [38] [^6] Genetic variance is composed of additive, dominance, and interaction variance. Here, we only consider additive variance, which is a common assumption in the field.. When selection occurs in each iteration, whether natural (e.g., faster animals surviving predators) or artificial (e.g., breeding livestock for higher milk production), it affects the distribution of the effective population that influences the next iteration. We consider directional selection with truncation, as shown in Figure 44(a), where a ratio of $r$ samples is selected by truncating $r_{1}$ from the left and $r_{2}$ from the right side of the distribution. Here, $r+r_{1}+r_{2}=1$ and larger phenotype values are preferred when $r_{1}>r_{2}$.
@@
where the mean phenotype of the next iteration $\mu_{t+1}$ is represented using the mean phenotype of selected individuals $\mu_{t}^{\prime}$, the mean phenotype of the entire population at current iteration $\mu_{t}$, and heritability $h_{t}^{2}$. Additionally, we assume the genetic variance for the next iteration is determined by the variance of selected individuals $\sigma_{G,t+1}^{2}=\sigma_{P,t}^{\prime 2}$. We prove the behaviors of mean and variance of phenotypes under this setting:
Theorem 1
@@
The proof of Theorem 1 is provided in Appendix D.1.
@@ 4.2 Explaining the Chain of Diffusion with Quantitative Trait Modeling
Just as phenotypes evolve through complex functions of hidden genotypes across multiple iterations in quantitative trait modeling, various continuous features of images evolve similarly during the Chain of Diffusion. A key discovery in our work is that the high-frequency power spectra are crucial phenotypes that evolve over iterations, with the CFG scales acting as different selection mechanisms. These power spectra are computed using 2D Fourier transforms, focusing on frequencies above a certain threshold to capture high-frequency components. Different CFG scales during generation correspond to different selection strategies: a high CFG selects individuals in the right tail of the distribution, favoring more detailed features with reduced diversity, while a low CFG selects from the left tail.
Figure 44(b) shows the evolution of power spectra in the Chain of Diffusion (in solid line) and the phenotype mean modeled with Equation  (9) (in dotted lines). The initial mean $\mu_{0}$ and genetic deviation $\sigma_{G,0}$ in the simulation are set to match those of the original image set at iteration 00. Three different configurations of ratios, $r_{1}$ and $r_{2}$, successfully model the power spectra distribution: CFG $7.5$ is modeled as selecting from upper $5$% of samples, CFG $2.5$ selects $50$% of the samples slightly favoring higher frequency, and CFG $1.0$ results from selecting the lower $30$%. We can see that this modeling can capture the evolution of power spectra very accurately over 6 iterations of Chain of Diffuson. Details about power spectra computation and simulation parameters are provided in Appendix D.2.2.
4.3 Statistical Genetics as a Unifying Lens to Model Collapse
Quantitative trait modeling, a tool borrowed from statistical genetics, serves as a solid theoretical framework for interpreting our experimental results. As genetic processes and the self-consuming training of generative models share a lot of similarities, we find that other concepts in mathematical genetics also have close ties to existing theoretical work on model collapse. For instance, the traditional Wright-Fisher model [19] describes how traits evolve from one generation to the next in a finite population. Due to randomness in the sampling process, the composition of traits within a finite population drifts over time and eventually collapses to a single phenotype. The Wright-Fisher model and its continuous variants [40] are exactly equivalent to the collapse behavior modeled with simple categorical or Gaussian distributions in [16, 3, 6, 22]. Given its widespread use in genetics, many extensions of the Wright-Fisher model remain an active area of research, such as including different mechanisms of selection [41, 42] or mutations [43]. Beyond the current work and existing theoretical studies on model collapse, we believe that statistical genetics offers a unified perspective to understand model collapse by reducing the complex dynamics of the self-consuming loop of generative model training to a small number of parameters, from which we can gain valuable insights.
@@ 5 Reusable Image Generation with ReDiFine
In the previous sections, we have discovered a significant role of CFG in model collapse and that an appropriate choice of CFG can effectively mitigate the collapse. However, the optimal CFG value that minimizes the collapse rate is different for each dataset (e.g., CFG 1.5 for CelebA, CFG 2.5 for Pokemon). There is no efficient way of finding an optimal CFG other than performing a grid search over a wide range of values and iteratively fine-tuning the diffusion model to evaluate the collapse rate for each configuration. In practical scenarios, it is unlikely that end users who are not machine learning experts will go through such a process during finetuning, just to prevent a potential model collapse that can happen in the next generation. This raises the question: How can we design a finetuning and generation strategy that is user-friendly and can effectively slow down model collapse?
To address this question, we again draw inspiration from the evolution process in nature. In biology, mutations naturally counteract genetic drift and preserve diversity. Furthermore, selection in nature is often a soft process rather than a hard truncation, as illustrated in Figure 44(a). This soft selection allows for the inclusion of outliers, thereby maintaining the overall genetic diversity. We connect these biological inspirations with two strategies: condition drop finetuning to introduce more randomness and CFG scheduling during generation to transform hard selection to a softer one [^8] We show a modified quantitative trait modeling result with these two modifications in Appendix D.2 and show that it captures the ReDiFine experimental results effectively. In this section, we focus on showing the image generation results with ReDiFine. We refer the curious readers to the appendix.. We propose Reusable Diffusion Finetuning (ReDiFine) method that incorporates these two ideas and show that it is highly effective at mitigating model collapse without the need for any finetuning.
@@
CFG Scheduling. We propose gradually reducing the CFG scale during diffusion steps to mitigate the impacts of overemphasizing the conditional score in later stages, which can lead to high-frequency degradation. Specifically, we exponentially decrease the CFG scale $s$ during $T$ diffusion steps as $s=s_{0}\times e^{-\alpha\times t/T}$, where $s_{0}$ is the initial CFG scale and $\alpha$ is the rate of exponential decay. This scheduling approach is consistent with findings by [44], which suggest that different diffusion steps contribute uniquely to the generation process. High CFG lets us capture high-level semantic information accurately during the initial diffusion steps while lower CFG in later steps can prevent generating unnecessary high-frequency details.
Figure 5: ReDiFine effectively mitigates image degradation in the Chain of Diffusion. ReDiFine successfully preserves the characteristics and features without further dataset-specific hyperparameter search. Artifacts observed in high-frequency degradation do not exist for all four datasets.
@@
Figure 6: ReDiFine effectively mitigates collapse-FID trade-off across all four datasets. While the optimal CFG scale varies for different datasets, ReDiFine consistently achieves low collapse rate and FID at the same time (lower is better). Note that the differences in FID are relatively smaller than those in collapse rate, supporting the necessity to evaluate collapse rate in the Chain of Diffusion.
@@
ReDeFine Results. We present the generated images and quantitative metrics for ReDiFine. We use the initial CFG scale $s_{0}=7.5$, diffusion steps $T=30$, decay rate $\alpha=2$, and condition drop probability $0.2$ for all of our experiments, except in the robustness comparison and ablation study. While ReDiFine can be further optimized through hyperparameter tuning (e.g., changing condition drop probability), our goal is to demonstrate that ReDiFine robustly mitigates model collapse with these default parameter settings for all datasets. In Figure 5, we can clearly see that ReDiFine significantly improves the image quality at later iterations compared to the baseline (Figure 11(b)) for all four datasets. In addition to the visual comparison, Figure 6 quantitatively shows the collapse-FID trade-off of ReDiFine. In all four datasets, ReDiFine shows substantially lower collapse rate (y-axis), compared to the baseline case (CFG=7.5). Furthermore, the performance of ReDiFine is close to the optimal Pareto curve spanned by different CFG scales, achieving similar performance as the optimal CFG values across all four datasets, demonstrating its effectiveness as a universal solution for mitigating model collapse. In contrast, using a fixed CFG scale that works well for one dataset often fails on others: CFG $2.5$ is optimal for Pokemon but performs poorly for CelebA-1k and Butterfly, and CFG $1.5$ is optimal for CelebA but performs poorly for Pokemon and Kumapi.
We also compare ReDiFine with real image mixing, which is suggested as a mitigation strategy in several previous papers [16, 3, 8, 23]. In Figure 7, we can observe that even infusing the dataset with 90% real images, it shows severe image degradation within six iterations. By visual comparison, we can see that Pikachu images in Figure 5 have less degradation than those with real image mixing. Furthermore, applying ReDiFine to these mixed datasets, as shown in Figure 7, significantly improves image quality compared to the baseline. This demonstrates that ReDiFine is an effective solution for various settings and can be used alongside other mitigation strategies such as real data mixing.
Figure 7: Comparison of ReDiFine and baseline (CFG scale 7.5) with training sets mixed with original and synthetic images from the previous iteration. While mixing real images as much as $90$% still results in model collapse for the baseline (as shown in the first row), the ability of ReDiFine to mitigate model collapse is even more highlighted by mixing real images (second row).
@@
Ablation study & Further analyses. We conducted an ablation study to understand the contributions of condition drop finetuning and CFG scheduling to the success of ReDiFine. We plot the results of using only condition drop and CFG scheduling in green triangles on Figure 6. In Pokemon, using only one of those strategies outperforms ReDiFine, but in all other datasets, using only one strategy shows higher collapse rate than ReDiFine. Especially in CelebA, using either one of them showed significantly worse performance than ReDiFine. These results suggest that combining both condition drop and CFG scheduling builds robustness to the method, making ReDiFine effective across all tested datasets. We conducted further analyses on ReDiFine, examining the distribution of latent features, the evolution of the norm of Diff, the power spectra using 2D Fourier transforms, and forensic fingerprints based on prior studies [45, 46]. Our analysis shows that ReDiFine effectively preserves the latent distribution and the norms of Diff over six iterations, with forensic fingerprints closely resembling those of the optimal CFG case. Detailed results are provided in Appendix G.
Reusable image generation as a responsible AI practice. In this section, we demonstrated that ReDiFine significantly slows the collapse rate without sacrificing image quality, similar to how fair classifiers reduce bias without compromising accuracy [47]. Adopting ReDiFine is a responsible, environment-conscious practice to prevent polluting the world with images that might look visually appealing, but totally unusable to improve future AI models.
@@ 6 Conclusion
The influx of AI-generated data into the world is inevitable and training sets that consist of synthetic data will be part of the AI development pipeline. In this paper, we empirically and theoretically studied the scenario of finetuning a model with its own generated data, where a gradual degradation called “model collapse” happens. We (1) identify the most impactful factor through comprehensive empirical investigations, (2) develop a novel theoretical perspective inspired by statistical genetics to explain model collapse, and (3) propose an effective mitigation strategy that generates reusable images for future training.
@@
We started this paper with a question: can current AI models learn from their own output and improve themselves? Our paper shows a glimpse that widely-used text-to-image models are not ready to improve from their own creation quite yet. While we presented one solution focused on generating reusable data, many open directions remain, such as developing algorithms that can distinguish between real and synthetic data and apply different learning techniques accordingly.
Supplementary Material
Appendix A Experimental setup
A.1 Hyperparameters

### v3

@@ 4 Quantitative Trait Models for Fully-Synthetic Training Loops
This section introduces a novel perspective to understand model collapse in generative models by drawing parallels between genetic biology and the Chain of Diffusion. Distinguishable iterations in the Chain of Diffusion—where each iteration is separable with no duplicated individual, and the current iteration originates from the previous one—mirror genetic processes involving successive iterations of parents and offspring. We use the term “iteration” instead of “generation” for genetic generation to avoid confusion with image generation.
We begin by explaining how “diversity collapse” of language models in self-consuming training loops [9, 11] could be explained by the classic Wright-Fisher model from statistical genetics. Then we introduce Quantitative Trait modeling and its underlying mathematical assumptions. Based on them, we derive a theorem showing that the mean trait value exhibits linear drift, while the variance stabilizes over time. This model can successfully capture the three key behaviors observed in Section 3: high-frequency degradation, low-frequency degradation, and optimal CFG scale. It suggests that different CFG scales in the Chain of Diffusion can be viewed as varying selection strategies, with power spectra corresponding to quantitative traits.
4.1 Connection to Statistical Genetics
The Wright-Fisher (WF) model [19] describes how allele (genetic variants) frequencies evolve in a population over iterations, assuming no selection or mutation. In each iteration, the entire population is replaced, with the next iteration formed through random sampling with replacement—a process known as genetic drift. When the population size is finite, the WF model predicts that one allele will eventually reach fixation, while others are lost. We provide further explanations in the Appendix E.1.
Language models (LMs) follow the same basic idea, with binary alleles replaced by a vocabulary of discrete words. LMs estimate word frequencies from a training corpus, and in self-consuming training loops, new texts are generated by the current LM and used to train the next. This acts analogously to genetic drift in large-scale, but finite, generations. Rare words may gradually disappear from the corpus—an effect that is irreversible. Recent works on diversity collapse in LMs [9, 11] can be interpreted as fixation by drift, in line with the WF model. In contrast, image generation operates in a near-continuous space of pixel combinations (e.g., $(256^{3})^{512\times 512}$ for $512\times 512$ RGB images), where no output is truly lost—unseen combinations can still be generated. In this setting, the process is typically modeled with Gaussian distributions [3, 6, 16, 23], and analogs of genetic drift and fixation emerge through the continuous variants of the WF model [45].
Given its widespread use in genetics, the WF model has been extended to incorporate selection [46, 47] or mutations [48], further enriching the analogy. These connections establish a strong conceptual correspondence between statistical genetics and model collapse—one that, to our knowledge, has been largely overlooked. Our work highlights this parallel and demonstrates how it can serve as a unifying theoretical perspective.
@@ 4.2 Quantitative Trait Modeling
Prior works can be viewed through the lens of genetic drift, where the persistence or disappearance of certain word tokens is driven largely by chance. In contrast, we adopt a Quantitative Trait (QT) modeling, incorporating a form of natural selection that reflects human preferences, to explain the evolution of feature distributions and the onset of model collapse. QT modeling in statistical genetics explores the evolution of quantitative phenotypes (e.g., height, weight, or color), which are influenced by many genetic and environmental factors. These traits are typically modeled as Gaussian-distributed variables and evolve across discrete iterations, where parent and offspring populations are distinguishable ($t$- and $t+1$-th iterations are clearly separable).
Let the distribution of phenotypes at the $t$-th iteration be denoted as $\mathcal{N}(\mu_{t},\sigma_{P,t}^{2})$ where $\mu_{t}$ and $\sigma_{P,t}^{2}$ are the mean and variance of quantitative phenotypes. The phenotypic variance $\sigma_{P,t}^{2}$ is the sum of the (additive) genetic variance $\sigma_{G,t}^{2}$ and the environmental variance $\sigma_{E}^{2}$, i.e., $\sigma_{P,t}^{2}=\sigma_{G,t}^{2}+\sigma_{E}^{2}$ [49] [^5] Genetic variance is composed of additive, dominance, and interaction variance. Here, we only consider additive variance, which is a common assumption in the field.. When selection occurs, whether natural (e.g., faster animals surviving predators) or artificial (e.g., breeding livestock for higher milk production), it affects the distribution of the effective population that influences the next iteration. We consider directional selection with truncation as shown in Figure 3(b), where $r$ ratio of samples is selected by truncating $r_{1}$ from the left and $r_{2}$ from the right side of the distribution. Here, $r+r_{1}+r_{2}=1$ and larger phenotypes are preferred when $r_{1}>r_{2}$.
@@
where the mean phenotype of the next iteration $\mu_{t+1}$ is represented using the mean phenotype of selected individuals $\mu_{t}^{\prime}$, the mean phenotype of the entire population $\mu_{t}$, and heritability $h_{t}^{2}$. We assume the genetic variance for the next iteration is determined by the variance of selected individuals $\sigma_{G,t+1}^{2}=\sigma_{P,t}^{\prime 2}$. We prove the behaviors of mean and variance of phenotypes under this setting:
Theorem 4.1.
@@
The proof of Theorem 4.1 is provided in Appendix E.2.
@@ 4.3 Explaining the Chain of Diffusion with Quantitative Trait Modeling
Just as phenotypes evolve through hidden genotypes in QT modeling, image features similarly evolve in the Chain of Diffusion. We identify high-frequency power spectra as a crucial phenotype, influenced by CFG scale as a selection mechanism. Using 2D Fourier transforms, we analyze high-frequency components above a certain threshold. Different CFG scales correspond to different selection strategies: a high CFG selects individuals in the right tail of the distribution, favoring more detailed features with reduced diversity, while a low CFG selects from the left tail.
Figure 3(c) compares the evolution of power spectra in the Chain of Diffusion (solid line) with the phenotype mean modeled by Theorem 4.1 (dotted lines). The simulation parameters, including initial mean $\mu_{0}$ and genetic deviation $\sigma_{G,0}$, are set to match the original image set at iteration 00. Three ratio configurations, $r_{1}$ and $r_{2}$, effectively model the power spectra distribution: CFG $7.5$ is modeled as selecting the top $5$%, CFG $2.5$ selects $50$% of the samples slightly favoring higher frequency, and CFG $1.0$ selected the lower $30$%. This modeling accurately captures power spectra evolution over six iterations of the Chain of Diffusion. More details can be found in Appendix E.3.2.
4.4 Further Discussions
Observable features of generated images are influenced by a complex interplay of factors like random noise, text signals, and various hyperparameters. The influence of CFG is thus multifaceted—analogous to how a single genotype gives rise to multiple phenotypes (e.g., red hair and freckles). Among many possible phenotypes affected by CFG, we identified one with a clear and consistent trend over iterations. It shows how the theoretical framework on mean-drift from QT modeling can interpret our empirical results, revealing that key ideas from statistical genetics naturally align with model collapse. Beyond the scope of our current findings, we believe that statistical genetics offers a unified perspective about model collapse by reducing complex dynamics of the self-consuming training loops to a few parameters, from which valuable insights can be drawn.
@@ 5 Reusable Image Generation with ReDiFine
In the previous sections, we have discovered a significant role of CFG in model collapse and that a good choice of CFG can mitigate the collapse while preserving the first iteration FID. However, the optimal CFG value differs for each dataset (e.g., 1.5 for CelebA, 2.5 for Pokemon), making iterative finetuning for evaluating the collapse rate on each configuration inevitable to find an optimal CFG. In practice, it is unlikely that non-expert end users will go through such a process to prevent a potential model collapse. This raises the question: How can we design a user-friendly finetuning and generation strategy that can slow down model collapse without CFG tuning?
Figure 4: ReDiFine significantly improves the baseline without extensive hyperparameter tuning in the Chain of Diffusion. It performs robustly across multiple datasets, resolving high-frequency degradation observed in the baseline with the default configuration.
To address this question, we again draw inspiration from the evolution process in nature where mutations naturally counteract genetic drift and preserve diversity. Furthermore, selection in nature is often a soft process rather than a hard truncation illustrated in Figure 3(b). This soft selection allows for the inclusion of outliers, thereby maintaining the overall genetic diversity. We connect these biological inspirations with two strategies: condition drop finetuning to include more randomness and CFG scheduling during generation to transform hard selection to a softer one [^7] We show a modified quantitative trait modeling result with these two modifications in Appendix E.3 and show that it captures the ReDiFine experimental results effectively.. We propose Reusable Diffusion Finetuning (ReDiFine) which integrates these two ideas and achieves a collapse rate similar to the optimal CFG, producing reusable images with minimal extra effort.
@@
Figure 5: ReDiFine performs comparably to the optimal CFG scale, outperforming the baseline (CFG=7.5) on collapse-FID trade-off across all four datasets. While the optimal CFG scale varies for different datasets, ReDiFine consistently achieves low collapse rate and FID at the same time (lower is better). Note that the differences in FID are relatively smaller than those in collapse rate, supporting the necessity to evaluate collapse rate in the Chain of Diffusion.
CFG Scheduling. We propose gradually reducing the CFG scale during diffusion steps to mitigate the impacts of overemphasizing the conditional score in later stages, which can lead to high-frequency degradation. Specifically, we exponentially decrease the CFG scale $s$ during $T$ diffusion steps as $s=s_{0}\times e^{-\alpha\times t/T}$, where $s_{0}$ is the initial CFG scale and $\alpha$ is the rate of exponential decay. This scheduling approach is consistent with findings by [51], which suggest that different diffusion steps contribute uniquely to the generation process. We note here that ReDiFine adopts exponential decrease of CFG scale for simplicity, but any advanced CFG scheduling methods [52, 53, 54] can be explored in future research.
@@
ReDeFine Results. We present the generated images and quantitative metrics for ReDiFine, showing its robust performance without hyperparameter tuning in all four datasets. We use the default initial CFG scale $s_{0}=7.5$, decay rate $\alpha=2$, and condition drop probability $0.2$, except in the robustness comparison and ablation study. Figure 4 shows that ReDiFine generates significantly better images compared to the baselines (CFG=7.5), and performs comparably to the optimal CFG scales. In addition to the visual comparison, Figure 5 quantitatively shows the collapse-FID trade-off of ReDiFine. In all four datasets, ReDiFine shows substantially lower collapse rate (y-axis), compared to the baseline case (CFG=7.5). Furthermore, the performance of ReDiFine is close to the optimal Pareto curve spanned by different CFG scales, achieving similar performance as the optimal CFG values, demonstrating its effectiveness as a universal and user-friendly solution. In contrast, using a fixed CFG scale that works well for one dataset often fails on others: CFG $2.5$ is optimal for Pokemon but performs poorly for CelebA-1k and Butterfly, and CFG $1.5$ is optimal for CelebA-1k but performs poorly for Pokemon and Kumapi. Further experimental results of ReDiFine for six additional iterations, cross-domain data, and iterative retraining can be found in Appendix F.
@@
Ablation study & Further analyses. We conducted an ablation study to understand the contributions of condition drop finetuning and CFG scheduling to ReDiFine. As shown in triangles on Figure 5, using only one of those strategies outperforms ReDiFine in Pokemon, but in all other datasets, using only one strategy shows higher collapse rate than ReDiFine. Especially in CelebA-1k, using either one of them showed significantly worse performance than ReDiFine. It suggests that combining condition drop and CFG scheduling builds robustness to the method, making ReDiFine effective across all tested datasets. We conducted further analyses on ReDiFine, examining the distribution of latent features, the evolution of the norm of Diff, the power spectra using 2D Fourier transforms, and forensic fingerprints based on prior studies [55, 56]. Our analysis shows that ReDiFine effectively preserves the latent distribution and the norms of Diff over six iterations, with forensic fingerprints closely resembling those of the optimal CFG case. Detailed results are provided in Appendix H.
@@ 6 Conclusion
The influx of AI-generated data into the world is inevitable and training sets that consist of synthetic data will be part of the AI development pipeline. In this paper, we empirically and theoretically studied the scenario of finetuning a generative model with its own generated data, where a gradual degradation called “model collapse” happens. We (1) identify the most impactful factor through comprehensive empirical investigations, (2) develop a novel theoretical perspective inspired by statistical genetics to explain model collapse, and (3) propose ReDiFine strategy for diffusion finetuning to slow down image quality degradation in model collapse.
@@
We started this paper with a question: can current AI models learn from their own output and improve themselves? Our paper shows a glimpse that widely-used text-to-image models are not ready to improve from their own creation quite yet. While we presented one solution focused on generating reusable data, many open directions remain, such as developing algorithms that can distinguish between real and synthetic data and apply different learning techniques accordingly. We further discuss the scopes and limitations of this work in Appendix A.
Appendix A Scope of Work, Limitations and Broader Impacts
A.1 Scope and limitations
In this paper, we provide three main contributions: significant observations from extensive experiments, a novel theoretical framework inspired by statistical genetics, and an effective mitigation strategy for model collapse. We detail the boundary and scope of our work in the literature to better understand our contributions.
First, our empirical observations are based on self-consuming training loops and finetuning of Stable Diffusion v1.5. Model collapse is a universal phenomenon observed in recent machine learning training, as high-quality generated contents are now available across multiple domains. We target diffusion finetuning with LoRA as it is one of the most practical interests in the field, where tons of finetuned models and images are being released to the Internet every day. While we already included the effects of mixing real images, iterative retraining, and domain accumulation to extend the setting from fully self-consuming training loops, we remain open to further investigations about full training, different finetuning methods, and architectures for future work.
Next, our theoretical framework focuses on introducing an interesting analogy between two fields to open a new area for discussion. We showed that QT modeling from statistical genetics provides a powerful tool to understand model collapse based on the structural analogy between self-consuming training loops and genetic processes. Further extensions based on advanced methods from statistical genetics would be an important future direction in investigating the model collapse of generative models. Specifically, finding connections between these advanced theoretical methods with self-consuming training loops augmented with real data or accumulated data would be an exciting future direction to extend this analogy.
Finally, the proposed ReDiFine is intended to be an efficient solution that can avoid multiple evaluations of the Chain of Diffusion while achieving performance comparable to that of the optimal CFG scale. It could perform comparably to the optimal CFG on four different datasets, which is a promising achievement as the first solution specific to diffusion finetuning. Investigating a more powerful method specific for diffusion finetuning would be a promising direction for future research, but we aim to introduce a simple method without additional real images or hyperparameter tuning that demonstrates the concept and importance of generating reusable images.
A.2 Broader Impacts
Generating reusable images for further training or finetuning holds the potential to significantly enhance the quality, stability, and robustness of future generative models. By promoting continuity across iterations, this approach can improve long-term performance and reduce degradation in self-consuming loops. However, as generative models become more powerful and self-reinforcing, there is an increasing risk that such advancements could be misused to replicate, mimic, or infringe upon copyrighted assets, licensed material, or the creative work of artists—especially when the boundary between generated and human-made content becomes less clear.
We strongly emphasize that any use of generated content for further training must be carried out with a deep understanding of the rights and intent associated with the original data. Respecting intellectual property, license agreements, and the creative labor of individuals is essential to maintaining ethical standards in machine learning research. While our work aims to improve the technical reliability of generative models, we advocate for responsible usage and encourage the community to develop frameworks and norms that protect creators while fostering innovation.
Appendix B Experimental setup
B.1 Hyperparameters